# Generalizing Medical Image Representations via Quaternion Wavelet Networks

### v3

@@ V-C Metrics
As objective evaluation metrics in the reconstruction task, we compute the structural similarity index (SSIM), the mean squared error (MSE), and the Fréchet inception distance (FID). While SSIM measures image degradation through structural information, FID estimates how the distributions of real and reconstructed images are far from each other.
For evaluating the results of the segmentation task instead, we compute the Dice similarity coefficient (DSC), the mean intersection over union (mIoU), and the mean absolute error (MAE). In order to perform a robust evaluation of the image modality translation task, we employ both a standard FID and an FID proposed in the original work [12]. The latter is a composition of images (16 slices) passed into a different model [53] and then evaluated with the Fréchet distance. We carefully follow the recipe in [12] for the computation of this task-specific metric.
VI Results and Discussion
In this section, we present the selected quaternion sub-bands we employ in our tests and the empirical evaluation we accomplish for validating our theoretical claims. We perform multiple runs with different seeds for each experiment and we report in the tables the average score and the standard deviation for each metric.
| Model | Params | SSIM | MSE | FID |
| CAE[35] | 0.9915 0.007 | 0.0006 0.0001 | 0.0046 0.0035 | |
| QUAVE-CAE | 0.9915 0.005 | 0.0006 0.0001 | 0.0036 0.0006 | |
| QCAE[35] | 0.9504 0.002 | 0.0015 0.0001 | 0.0293 0.0024 | |
| QUAVE-QCAE | 0.9842 0.002 | 0.0010 0.0001 | 0.0126 0.0034 |
| Model | Params | DSC | mIoU | MAE |
| U-Net [51] | M | 0.74743 0.05954 | 0.65463 0.01655 | 0.06525 0.00300 |
| QUAVE-U-Net | 0.77257 0.00252 | 0.63937 0.00938 | 0.06661 0.00153 | |
| U-Net++ [52] | M | 0.77960 0.00463 | 0.63790 0.01473 | 0.06635 0.00155 |
| QUAVE-U-Net++ | 0.79813 0.01066 | 0.67290 0.01094 | 0.06142 0.00244 | |
| QU-Net[35] | M | 0.76680 0.01159 | 0.62933 0.02033 | 0.06887 0.00249 |
| QUAVE-QU-Net | 0.77280 0.01022 | 0.63820 0.01585 | 0.06632 0.00273 | |
| QU-Net++[35] | M | 0.78570 0.00776 | 0.64963 0.02230 | 0.06558 0.34956 |
| QUAVE-QU-Net++ | 0.79117 0.00196 | 0.65257 0.00386 | 0.06355 0.00029 |
| Model | FID[12] | FID | IS | DSC | S-Score | mIoU | MAE |
| TarGAN [12] | 0.0912 0.0002 | 0.7009 0.1181 | 3.1879 0.2172 | 87.53 0.6474 | 76.23 9.5988 | 0.7812 0.0090 | 0.0489 0.0007 |
| QUAVE-TarGAN | 0.0893 0.0025 | 0.6295 0.0233 | 3.3452 0.0185 | 90.53 4.3608 | 84.41 4.2848 | 0.8320 0.0676 | 0.0487 0.0004 |
| QTarGAN | 0.0881 0.0021 | 0.6513 0.0144 | 2.7842 0.2628 | 83.56 1.1758 | 76.85 1.2373 | 0.7246 0.0158 | 0.0492 0.0011 |
| QUAVE-QTarGAN | 0.0902 0.0024 | 0.6253 0.0860 | 3.0228 0.0524 | 83.02 0.7664 | 78.90 1.1821 | 0.7360 0.0096 | 0.0488 0.0005 |
VI-A Evaluation
VI-A1 Reconstruction on IXI dataset
Table I shows the average and the standard deviation over different runs of the objective metrics results for the IXI dataset. The models that employ the proposed preprocessing (QUAVE) enhance the scores with both real and quaternion-valued models. These results prove that QUAVE can boost model learning without requiring modification or manipulation of models or any kind of data augmentation. Indeed, features extracted with the QWT capture salient characteristics of the image that are crucial for proper reconstruction.
VI-A2 Segmentation on Kvasir-SEG dataset
Table II reports instead objective metrics results for the segmentation task on the Kvasir-SEG dataset with RGB images. It is worth noting that the Kvasir-SEG dataset comprises RGB images, thus the real-valued U-Net is trained with three-channel color samples. Therefore, we prove that our technique, QUAVE, increases model performance also with respect to color images. Hence, QUAVE can be generalized to grayscale and color images as well, improving the accuracy of involved networks.
@@ VI-A2 Segmentation on Kvasir-SEG dataset
CHAOS |
QUAVE-TarGAN | DWT | - | - | 0.092 0.002 | 0.684 0.040 | 0.888 0.201 | 0.802 0.031 | 0.048 0.001 |
@@ VI-A3 Image modality translation on CHAOS dataset
Table III reports objective metrics results for the image modality translation task on the CHAOS dataset. As in previous experiments, we apply the proposed QUAVE to data for extraction, to both the original TarGAN and a quaternion version referred to as QTargan. We can observe that both the DWT and QUAVE improve the performance with respect to the original network. The FID[12] proposed in the original TarGAN work tends to reward the images that best translate meaningful features in medical images, and our approach QUAVE beats the original processing model. Even if we consider the classical FID, QUAVE performs best among the original work. We also show this effect on a visual comparison among the predicted translation in Fig. 5 on randomly chosen samples. We display the superiority of the translation of QUAVE using red boxes. The boxes help to indicate the artifacts in the translation committed by the other methods. To conclude, we show that the data-agnostic QUAVE can be easily generalized to multimodality datasets, improving the performance of large-scale models too.
@@ VI-B Ablation study
We adopt the QWT for our framework, consequently, we show an ablation study instead of using the DWT in QUAVE. This is possible with only a slight change in the output size of the final layers of QUAVE. From the scores in Tab. IV, we can observe that both the DWT and QUAVE approaches have a high impact on model performance, improving scores with respect to baselines. However, the best scores are, again, achieved by the models with the QUAVE method with QWT wavelet, proving its effectiveness in most of all the different tasks.
@@ VII Conclusion
In this paper, we have introduced QUAVE a generalizable, shift-invariant, and data-agnostic framework, for exploiting salient features in medical images. The method is based on the quaternion wavelet transform and on a novel technique for extracting informative sub-bands within the Quaternion Wavelet Transform for neural models which improves networks’ performance in a variety of tasks, while improving quaternion and hypercomplex models’ generalizability. On a thorough experimental evaluation with different models in three distinct tasks, we have demonstrated the effectiveness and the generalization ability of our approach, which is able to improve performance according to a variety of metrics in each task we consider. Moreover, we have shown the flexibility of the proposed methods that can be involved in every medical and non-medical dataset, whether with grayscale, RGB images, or comprising different modalities.

### v5

@@ V-C Metrics
As objective evaluation metrics in the reconstruction task, we compute the structural similarity index (SSIM), the mean squared error (MSE), and the Fréchet inception distance (FID). While SSIM measures image degradation through structural information, FID estimates how the distributions of real and reconstructed images are far from each other. Indeed, FID is traditionally used in generative tasks where paired metrics are unavailable. In reconstruction tasks, its ability to capture differences in the distribution of high-level image features can offer complementary insights into reconstruction fidelity, particularly in terms of global structure and perceptual coherence. By including it alongside SSIM and MSE, we ensure a more comprehensive evaluation of our method ability to reconstruct high-quality images while maintaining both local and global features. For evaluating the results of the segmentation task instead, we compute the Dice similarity coefficient (DSC), the mean intersection over union (mIoU), and the mean absolute error (MAE). In order to perform a robust evaluation of the image modality translation task, we employ both a standard FID and an FID proposed in the original work [12]. The standard FID metric measures the distance between the distributions of feature vectors extracted from real and generated images using a pretrained network, commonly the Inception v3 model [58], which is trained on generic (non-medical) images. In the classic implementation, feature vectors from the last pooling layer of the pretrained Inception model are used to calculate the mean and covariance for both the real and generated images. Instead, the FID implemented by TarGAN [12] incorporates specific adaptations tailored for the task of translating medical images. To compute the FID, the feature vectors are extracted from the last layer of the encoder of a pre-trained Models Genesis [59] network and not the Inception v3 model like in the classic FID. Since Models Genesis [59] is based on 3D medical volume data, it is crucial that the input to this network also has a 3D context. In our image translation task, only 2D translations are performed. To provide the required 3D context for feature extraction, we form 3D volume patches by concatenating every 16 consecutive axial slices, following the receipt in [59]. The choice of 16 is based on a balance between capturing sufficient volumetric context and maintaining computational efficiency. Importantly, we ensure that the slices are anatomically consistent, i.e. they are consecutive so that the resulting 3D patch reflects the true spatial continuity of the anatomy. This ensures that the features extracted by Models Genesis are meaningful and appropriate for FID computation. The average FID score across all modality pairs (e.g., CT T1w, T2w T1w) is reported. Performance differences may exist between individual translation pairs, for example, translating from CT to T1w typically yields a higher FID score than translating from T1w to CT, owing to inherent contrast differences between these modalities. However, our analysis shows that despite these individual variations, the average FID remains a robust indicator of the overall quality and consistency of the translation process. In this context, the averaged metric effectively captures the general performance of our approach across all modality pairs, demonstrating that our method handles varying degrees of translation difficulty competently.
| Model | FID[12] | FID | IS |
| TarGAN [12] | |||
| QUAVE-TarGAN | (+2.1%) | (+10.1%) | (+4.9%) |
| QTarGAN | |||
| QUAVE-QTarGAN | () | (+3.9%) | (+8.5%) |
@@
| CHAOS | QUAVE-TarGAN | DWT | - | - | 0.092 0.002 | 0.684 0.040 | 0.888 0.201 | 0.802 0.031 | 0.048 0.001 |
@@
VI Results and Discussion
In this section, we present the selected quaternion sub-bands we employ in our tests and the empirical evaluation we accomplish for validating our theoretical claims. We perform multiple runs with different seeds for each experiment and we report in the tables the average score and the standard deviation for each metric. We initialized the neural networks weights (both the 2D convolutional and the linear layers) using the well-known Kaiming and He initialization with a uniform distribution and the biases to zero.
VI-A Evaluation
VI-A1 Reconstruction on IXI dataset
Table II shows the average and the standard deviation over different runs of the objective metrics results for the IXI dataset. The models that employ the proposed preprocessing (QUAVE) enhance the scores with both real and quaternion-valued models. These results prove that QUAVE can boost model learning without requiring modification or manipulation of models or any kind of data augmentation. Indeed, features extracted with the QWT capture salient characteristics of the image that are crucial for proper reconstruction.
VI-A2 Segmentation on Kvasir-SEG dataset
Table III reports instead objective metrics results for the segmentation task on the Kvasir-SEG dataset with RGB images. It is worth noting that the Kvasir-SEG dataset comprises RGB images, thus the real-valued U-Net is trained with three-channel color samples. Therefore, we prove that our technique, QUAVE, increases model performance also with respect to color images. Hence, QUAVE can be generalized to grayscale and color images as well, improving the accuracy of involved networks.
@@ VI-A3 Image modality translation on CHAOS dataset
Table IV and Table I report objective metrics results for the image modality translation task on the CHAOS dataset with both image generation and segmentation metrics. The segmentation results presented in Table I provide insight into the quality of structural preservation in the translated images. This aspect is inherently linked to the TarGAN architecture and methodology.
As in previous experiments, we apply the proposed QUAVE to data for extraction, to both the original TarGAN[12] and a quaternion version referred to as QTargan. To further stabilize GAN training, we also added to the original network the Spectral Normalization to limit the discriminator sensitivity and ensure a more stable training process [60]. We can observe that both the DWT and QUAVE improve the performance with respect to the original network. The FID[12] proposed in the original TarGAN work tends to reward the images that best translate meaningful features in medical images, and our approach QUAVE beats the original processing model. Even if we consider the classical FID, QUAVE performs best among the original work. We also show this effect on a visual comparison among the predicted translation in Fig. 6 on randomly chosen samples. We display the superiority of the translation of QUAVE using red boxes. The boxes help to indicate the artifacts in the translation committed by the other methods. To conclude, we show that the data-agnostic QUAVE can be easily generalized to multimodality datasets, improving the performance of large-scale models too.
@@ VI-B Ablation study
We adopt the QWT for our framework, consequently, we show an ablation study instead of using the DWT in QUAVE. This is possible with only a slight change in the output size of the final layers of QUAVE. From the scores in Table V, we can observe that both the DWT and QUAVE approaches have a high impact on model performance, improving scores with respect to baselines. However, the best scores are, again, achieved by the models with the QUAVE method with QWT wavelet, proving its effectiveness in most of all the different tasks.
@@ VII Conclusion
In this paper, we have introduced QUAVE a generalizable, shift-invariant, and data-agnostic framework, for exploiting salient features in medical images. The method is based on the quaternion wavelet transform and on a novel technique for extracting informative sub-bands within the Quaternion Wavelet Transform for neural models which improves networks’ performance in a variety of tasks, while improving quaternion and hypercomplex models’ generalizability. On a thorough experimental evaluation with different models in three distinct tasks, we have demonstrated the effectiveness and the generalization ability of our approach, which is able to improve performance according to a variety of metrics in each task we consider. Moreover, we have shown the flexibility of the proposed methods that can be involved in every medical and non-medical dataset, whether with grayscale, RGB images, or comprising different modalities. As future work, QUAVE may be defined to directly operate with 3D volumes of medical imaging inputs and outputs, so as to enable 3D downstream models in both real- and quaternion-valued domains.